OpenStudy (anonymous):

Christina Hercher borrowed $50,000 on a 90 day, eight percent note. Christina paid $3,000 toward the note on day 40. On day 60 she paid an additional $4,000. Using the U.S. Rule, Christina's adjusted balance after the first payment is

OpenStudy (anonymous):

50,000x0.08=4000+3000=7000+4000=$11000

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

depends on where is first payment is, assuming 1 payment is 90 days

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

that might be wrong i'm still calculating.

OpenStudy (anonymous):

54000 including interest, minus 7000 payments is 47000